OpenStudy (anonymous):

Find the sum of the following infinite geometric series, if it exists. 1/2 + (-1/4) + 1/8 + (-1/16) +...

OpenStudy (welshfella):

Sum to infinity = a1 /( 1 - r)

OpenStudy (welshfella):

where a1 = first term and r = common ratio

OpenStudy (anonymous):

so far i have r=-1/2

OpenStudy (welshfella):

thats correct

OpenStudy (anonymous):

does S=1/4?

OpenStudy (anonymous):

i'll put my work on a whiteboard

OpenStudy (anonymous):